OpenStudy (anonymous):

MEDALS FOR THE BEST ANSWER! PLEASE HELP! Solve the quantity x minus 5, end quantity, over x, minus 1, equals the quantity x plus 3, end quantity, over x.

OpenStudy (anonymous):

\[\frac{x-5}{x-1}=\frac{x+3}{x}\] is my guess

OpenStudy (anonymous):

"cross multiply" and start with \[x(x-5)=(x-1)(x+3)\] and solve the resulting quadratic equation

OpenStudy (anonymous):

which isn't *really* a quadratic at all...

OpenStudy (anonymous):

good point

OpenStudy (anonymous):

ok thank u!! @mtbender74 @satellite73

OpenStudy (anonymous):

yw